Directions: find (f-g)(x) using the two functions given in each number.



1.f(x) =3 x


g(x) =-4x+1



2.f(x) =2x


g(x)= 4x2+2x-2



3.f(x) =5x+1


g(x) =3x



4.f(x) =5x+2


g(x)=3x-5



5.f(x) =3x


g(x) =4x²+5x-4


1
Expert's answer
2022-04-05T15:50:12-0400

"1:\\\left( f-g \right) \left( x \right) =3x-\left( -4x+1 \right) =7x-1\\2:\\\left( f-g \right) \left( x \right) =2x-\left( 4x^2+2x-2 \right) =-4x^2+2\\3:\\\left( f-g \right) \left( x \right) =5x+1-3x=2x+1\\4:\\\left( f-g \right) \left( x \right) =5x+2-\left( 3x-5 \right) =2x+7\\5:\\\left( f-g \right) \left( x \right) =3x-\left( 4x^2+5x-4 \right) =-4x^2-2x+4"